OpenStudy (anonymous):

If there are, on average, 365.25 days in one year, how many seconds are in one year?

OpenStudy (anonymous):

whats up with these easy conversion questions bro?

OpenStudy (anonymous):

\[365.25\frac{days}{year}\times24\frac{hours}{day}\times60\frac{minutes}{hour}\times60\frac{seconds}{minute}=\cdots\frac{seconds}{year}\]

OpenStudy (anonymous):

does it matter? i need these answers cause i cant seem to get them.

OpenStudy (anonymous):

\[365.25\frac{days}{year}\times24\frac{hours}{day}\times60\frac{minutes}{hour}\times60\frac{seconds}{minute}=\cdots\frac{seconds}{year}\]

OpenStudy (anonymous):

ohh bro it doesnt just because they are easy for me doesnt me their easy for you too goodluck